### v2

# CompilerDream: Learning a Compiler World Model for General Code Optimization

Abstract
Effective code optimization in compilers is crucial for computer and software engineering. The success of these optimizations primarily depends on the selection and ordering of the optimization passes applied to the code. While most compilers rely on a fixed sequence of optimization passes, current methods to find the optimal sequence either employ impractically slow search algorithms or learning methods that struggle to generalize to code unseen during training. We introduce CompilerDream, a model-based reinforcement learning approach to general code optimization. CompilerDream comprises a compiler world model that accurately simulates the intrinsic properties of optimization passes and an agent trained on this model to produce effective optimization strategies. By training on a large-scale program dataset, CompilerDream is equipped to serve as a general code optimizer across various application scenarios and source-code languages. Our extensive experiments first highlight CompilerDream’s strong optimization capabilities for autotuning, where it leads the CompilerGym leaderboard. More importantly, the zero-shot generalization ability of large-scale trained compiler world model and agent, excels across diverse datasets, surpassing LLVM’s built-in optimizations and other state-of-the-art methods in both settings of value prediction and end-to-end code optimization.
1 Introduction
Code optimization plays an important role in realizing the full potential of software and hardware. Developers desire a universal solution to transform input programs into semantically equivalent but more efficient versions without manual effort. To achieve this, compilers usually employ a front-end that translates the source code of any language into an intermediate representation (IR), a back-end that converts IR to the binary code optimizer, and a middle-end optimizer that performs language- and platform-agnostic optimizations over the IR of the original code (Figure 1). The optimizer is typically implemented as a series of passes applying transforms on the code. The performance of the optimization primarily depends on the selection and order of the optimization passes. Standard compilers use a few fixed sets of optimization sequences to enhance specific aspects of program performance. For example, -O1, -O2, and -O3 improve for execution speed, and -Os and -Oz for reducing program size. Obviously, given the vast diversity of programs and platforms, these off-the-shelf optimization strategies predefined by compiler experts are not optimal for most circumstances, and we can significantly improve the performance of a specific program if we can find a better optimization pass sequence compared to that provided by compilers (Triantafyllis et al. 2003; Georgiou et al. 2018).
Figure 1: CompilerDream agent performs code optimization using a learned policy in a compiler environment. Guided by a learned policy, at each step, CompilerDream analyzes the Intermediate Representation (IR) of the program, selects to apply an optimization pass, and receives a reward based on the improvement in optimization metrics. Through sequential interactions, the agent aims to maximize cumulative rewards and enhance the final IR’s performance.
Automatically optimizing the pass sequence is thus crucial in improving the compiler optimization performance. To be practical, such an algorithm must produce a satisfactory pass sequence within a reasonable time and handle a wide variety of programs. However, current research often fails to meet these requirements simultaneously. Search-based methods (Bodin et al. 1998) ensure near-optimality but require thousands of compilations for each new program, making it too time-consuming for regular use. In contrast, machine learning methods can either predict the optimization sequences directly or estimate the outcomes of applying an optimization sequence to guide a search, thereby avoiding the time-consuming process of interacting with the compiler to validate the optimization outcome. However, in addition to potentially sacrificing some optimality compared to searching, these learning-based methods face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples.
A range of machine learning applications (Devlin et al. 2019; Brown et al. 2020; Radford et al. 2021; Kirillov et al. 2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance. Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al. 2022). However, the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al. 2020; Shahzad et al. 2022) or from relatively small training sets comprising only a few hundred programs (Mammadli, Jannesari, and Wolf 2020; Jain et al. 2022). Additionally, the predominant use of model-free reinforcement learning methods with limited capacity may also hinder generalization.
In this study, we focus on the LLVM (Lattner and Adve 2004) phase ordering problem, a longstanding challenge for compiler research, and propose CompilerDream, a world model-based reinforcement learning approach capable of generating an optimization pass sequence tailored to any particular input program. Our model-based method (Hafner et al. 2023) not only learns an accurate predictive world model of compilers to simulate real compiler executions, whose number can be greatly reduced, but also benefits generalization by learning representations that better capture the structure of the compiler state transitions (Anand et al. 2022). Trained on a carefully curated large-scale dataset of natural programs, the world model accurately captures the dynamics of the compiler optimization process, enabling the optimization agent to generalize to unseen programs and predict superior optimization sequences.
We demonstrate the effectiveness of CompilerDream on a range of domains (Cummins et al. 2022), from benchmark suites encompassing fundamental algorithms to production-level open-source programs, including object files from C++ TensorFlow (Abadi et al. 2016) and OpenCV (Culjak et al. 2012) library. Even without considering generalization, CompilerDream’s strong optimization capabilities place it at the top of the CompilerGym leaderboard for autotuning. For general code optimization, the large-scale trained world models of CompilerDream can accurately predict the outcomes of pass sequences on unseen programs, a setting known as value prediction. Beyond selecting optimization sequences based on value prediction, the agent can also directly produce optimization sequences in a single trial, yielding more efficient code size reduction compared to both the -Oz flag and state-of-the-art methods.
The main contributions of this work are three-fold:
  • We introduce CompilerDream, the first model-based reinforcement learning approach for code optimization.
  • We first leverage a large-scale dataset, CodeContests, to achieve strong generalization on unseen programs.
  • Extensive experiments spanning autotuning, value prediction, and end-to-end code optimization, demonstrate the superior performance of CompilerDream.
2 Related Work
One of the key challenges for compilation is to determine which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order. This requires effectively searching and evaluating a massive number of possible options, known as iterative compilation (Bodin et al. 1998) or autotuning (Datta et al. 2008). However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler strategy. This limitation underscores the importance of integrating machine learning techniques.
Supervised Learning
Pioneering work has delved into supervised machine learning, adopting two main approaches (Leather and Cummins 2020). The first approach requires an extensive search on each training programs to identify the most effective optimization sequence, which then serves as the data labels. An early example (Calder et al. 1997) used a neural network for branch prediction, and one more well-known work is MilepostGCC (Fursin et al. 2008), a practical attempt to integrate machine learning into a production compiler, GCC. It employs models trained on a large dataset of programs distributed over the Internet. The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables evaluation of a range of possible options without the need to compile and profile each one (Stephenson et al. 2003; Luk, Hong, and Kim 2009).
Reinforcement Learning
Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the requirement for collecting optimal labeled data (Kulkarni and Cavazos 2012). This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al. 2021), loop transformation (Haj-Ali et al. 2020; Brauckmann, Goens, and Castrillón 2021), and graph partitioning (Mirhoseini et al. 2017). Several works relevant to us have explored the full optimization pipeline, i.e. the LLVM phase ordering problem, including AutoPhase (Haj-Ali et al. 2020), CORL (Mammadli, Jannesari, and Wolf 2020), and POSET-RL (Jain et al. 2022).
Model-based Reinforcement Learning
While existing work on reinforcement learning for code optimization predominantly focuses on model-free RL methods, our work pioneers the use of an advanced model-based RL approach. Model-based RL (Sutton and Barto 2018) can offer advantages in terms of both sample efficiency and generalization. Executing and profiling extensive optimization sequences, especially for the runtime metric, can be time-consuming. This is further compounded when constructing complex observations, such as control-data flow graphs (Brauckmann et al. 2020; Cummins et al. 2021). Model-based RL addresses these challenges by learning a world model that approximates state transitions and reward signals of the compiler. This is related in spirit to the aforementioned approach of supervised performance models but goes further by also learning a policy through simulating trajectories based on model predictions. This approach thus improves sample efficiency not only by eliminating the necessity for a guided search process during deployment but also by reducing trial-and-error interactions with the real compiler environment during training. Moreover, as the policy can share the representation with the world model, model learning can act as an auxiliary task and thus aid in learning representations that better capture the structure of the environment and manifest in better generalization of the policy (Anand et al. 2022; Mazoure et al. 2022; Ma et al. 2024).
Figure 2: Design overview of CompilerDream. (a) In the model-based training loop, the CompilerDream agent interacts with the compiler environment across diverse training programs, learns a world model from historical experience, and updates its policy efficiently through model simulations. (b) The trained agent can be deployed and is capable of zero-shot generalization to unseen programs, delivering effective optimizations. (c) The compiler environment is set up with the input program’s IR, $s_{0}$. At each step, upon receiving an optimization pass $a_{t}$, the environment executes the pass internally, resulting in a transformed IR $s_{t+1}$, and provides the agent with relevant IR features $o_{t+1}$ and immediate rewards $r_{t+1}$ based on optimization metric improvements. (d) The world model simulates the compiler environment, initiated with an observation, maintaining its internal neural states, and predicting future observations and rewards in response to input actions. (e) The policy shares with the world model a generalizable representation that captures the environment’s structure.
3 Method
In this section, we first present the mathematical formulation of the code optimization problem (Section 3.1). Then we introduce the CompilerDream approach for both compiler world model learning (Section 3.2) and optimization agent learning (Section 3.3). Finally, we elaborate on considerations of curating large-scale training dataset (Section 3.4).
3.1 Phase Ordering Decision Process
As illustrated in Figure 1, one key problem of compiler optimization is to find the optimal sequence of optimization passes for a given program, which is also referred to as the phase ordering problem. It can be naturally formulated as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$ (Sutton and Barto 2018). The state space $\mathcal{S}$ covers all possible Intermediate Representations (IRs), the action space $\mathcal{A}$ comprises individual compiler optimization passes, and the reward function $r$ is defined by the metric being optimized. The transition dynamics $p:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ represents the outcome of applied IR transformations. The initial state distribution $\mu\in\Delta(\mathcal{S})$ captures all IRs of interest, which can be approximated via uniform sampling from the training dataset. The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space. Complex observation spaces offer comprehensive program information, while expert-designed features (Haj-Ali et al. 2020) embed problem-specific knowledge, potentially enhancing the optimization policy’s generalization by eliminating irrelevant details.
A code optimization agent guided by policy $\pi$ determines the optimization sequence of an input program through a series of interactions with the compiler environment. At each time step $t=0,1,2,\dots$, the agent applies an action $a_{t}$ to current IR based on observation $o_{t}$ and its policy $\pi(a_{t}\mid o_{\leq t})$, which transforms the current state $s_{t}$ to the next state $s_{t+1}=p(s_{t},a_{t})$ and will receive a reward $r_{t}$. The goal of the agent is to learn an optimization policy $\pi$ that effectively maximizes the cumulative rewards $G=\sum_{t}r_{t}$, thereby optimizing the performance metrics of the final IR.
3.2 Learning a Compiler World Model
We develop our CompilerDream approach by adapting an advanced model-based RL method, Dreamer (Hafner et al. 2023), as depicted in Figure 2. This approach involves learning a predictive world model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment parameterized by $\theta$ that approximates the underlying transition dynamics $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and reward function $r(o_{\leq t},a_{\leq t})$.
The compiler world model simulating the compiler environment is formulated with the following four components:
Representation model: $z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})$ (1)
Transition model: $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})$
Observation decoder: $\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t})$
Reward decoder: $\hat{r}_{t}\sim p_{\theta}(\hat{r}_{t}\mid z_{t}).$
The representation model estimates a neural compiler state $z_{t}$ from the current observation $o_{t}$ of real compiler state, the previous state $z_{t-1}$ and the previous optimization action $a_{t-1}$. A representation loss $\mathcal{L}_{\mathrm{repr}}$ is used to train the neural compiler states to accurately reconstruct the observation and reward by two decoders:
$\mathcal{L}_{\mathrm{repr}}(\theta)\doteq-\ln p_{\theta}(o_{t}|z_{t})-\ln p_{\theta}(r_{t}|z_{t}).$ (2)
The transition model captures the dynamics of the compiler world model, predicting future neural state ${\hat{z}}_{t}$ directly from $z_{t-1}$ and $a_{t-1}$. A prediction loss minimizes the difference between the estimated state $z_{t}$ and the predicted state ${\hat{z}}_{t}$, simultaneously enhancing the transition model’s accuracy in predicting future states and making the state produced by the representation model easier to predict:
$\mathcal{L}_{\mathrm{pred}}(\theta)\doteq\mathrm{KL}\left[q_{\theta}(z_{t}|z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}|z_{t-1},a_{t-1})\right].$ (3)
The overall models are jointly learned by minimizing the sum of the representation loss and the prediction loss.
Compiler simulation
We can simulate the behavior of the real compiler using our trained compiler world model. Specifically, a simulated compiler optimization trajectory $\{\hat{z}_{\tau},\hat{a}_{\tau},\hat{r}_{\tau},\hat{o}_{\tau}\}$ with horizon $H$ can be generated by the interactions between the world model and an optimization agent: starting at a neural state $\hat{z}_{t}\sim q_{\theta}(z_{t}|o_{t})$, at each step $\tau=t,t+1,t+2,\dots$, the agent takes an action $\hat{a}_{\tau}\sim\pi_{\psi}\left(\hat{a}_{\tau}|\hat{z}_{\tau}\right)$, and transits to the next latent state $\hat{z}_{\tau+1}\sim p_{\theta}(\hat{z}_{\tau+1}|z_{\tau},a_{\tau})$ with a reward $\hat{r}_{\tau+1}\sim p_{\theta}(\hat{r}_{\tau+1}|\hat{z}_{\tau+1})$. Predicted real compiler state $\hat{o}_{\tau+1}$ can be optionally reconstructed by the observation decoder.
Reward smoothing
During an optimization sequence, not all passes can effectively transform the code, and the improvements gradually saturate. We employ a reward smoothing technique (Lee, Abbeel, and Lee 2024) to mitigate the sparsity and the long tail distribution of rewards during an episode. This is achieved through the following:
$r_{t}^{\prime}\leftarrow\alpha r_{t-1}^{\prime}+(1-\alpha)r_{t},\;\;t=1,2,\dots$ (4)
with $\alpha\in[0,1)$. Consequently, we train a reward decoder $p_{\theta}(\hat{r}_{t}^{\prime}\mid z_{t})$ to predict the smoothed rewards.
3.3 Learning a Code Optimization Agent
We then train a code optimization agent on top of the world model, which consists of an actor and a critic neural networks parameterized on top of the neural compiler state:
[…]
The critic evaluates the cumulative optimization return that the actor can achieve starting from a given neural state, also known as value or return, denoted as $v_{\xi}(\hat{z}_{t})$. This is trained by minimizing the difference between the predicted value $v_{\xi}(\hat{z}_{t})$ and the Monte-Carlo return or more advanced bootstrapped $\lambda$-return (Sutton and Barto 2018):
$\mathcal{L}_{\text{critic}}(\xi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}-\log v_{\xi}(V_{\tau}^{\lambda}\mid\hat{z}_{\tau})\right],$ (6)
$V_{\tau}^{\lambda}\doteq\hat{r}_{\tau}+\gamma\begin{cases}(1-\lambda)v_{\xi}(\hat{z}_{\tau+1})+\lambda V_{\tau+1}^{\lambda}&\text{if}\ \tau<t+H\\ v_{\xi}(\hat{z}_{\tau+1})&\text{if}\ \tau=t+H.\end{cases}$
The actor produces a policy $\pi_{\psi}$ that predicts the best optimization pass to choose, represented as an action distribution, which is trained to maximize the simulated return through the REINFORCE policy gradient (Williams 1992):
$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}^{\lambda}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}|\hat{z}_{\tau})$
$-\eta\,\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\Big{)}\Big{]},$ (7)
where $\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]$ is an entropy regularization which encourages exploration, and $\eta$ is a hyperparameter that adjusts the regularization strength.
Both the actor and critic are completely trained on world model-simulated compiler optimization trajectories.
3.4 Data Curation
To facilitate that our CompilerDream method can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset. First, the dataset should reflect naturalness. During our preliminary experiments, we found that those generated code datasets like Csmith (Yang et al. 2011) and llvm-stress (Lattner and Adve 2004) provide no benefits or even hurt the generalization to real-world scenarios. Second, the dataset should be large to prevent the agent from overfitting to a small number of codes and failing to generalize. Last, the code must exhibit high quality from an optimization perspective. We need data with complex algorithmic logic and potential for optimizations to help the world model and the agent better understand the problem’s intricacies. Datasets like AnghaBench (Da Silva et al. 2021) which consist of millions of human-written codes collected from GitHub, are often too simple to allow for significant optimization improvements.
Therefore, we choose to construct our training datasets on top of the CodeContests dataset released with AlphaCode (Li et al. 2022), which consists of over 13,000 problems of coding competition, and each problem, on average, has hundreds of solutions in multiple languages. We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as our training data, and sample one solution for each of 100 test problems as our validation data.
4 Experiments
We evaluate CompilerDream in three different settings and compare its performance with prior state-of-the-art methods. We first demonstrate that, without considering generalization, CompilerDream excels as a powerful autotuning method, leading the CompilerGym leaderboard (Section 4.2). Then we show that CompilerDream can learn a general compiler world model capable of accurately estimating and selecting pass sequences (Section 4.3) and finally learn a general optimization agent that directly generates a pass sequence tailored to any specific input program. (Section 4.4)
4.1 Evaluation
Our experiments focus on code size reduction, which benefits applications targeting low-resource hardware such as embedded systems. This focus is driven by the practical advantages of code size as a metric: it is cost-effective to construct extensive compilable training and test datasets and to evaluate the optimization performance for code size.
Metric
To be more robust to outliners, we evaluate the code size optimization results by the geometric mean of IR instruction count reduction of program $s$ in a dataset $\mathcal{D}$:
$R_{\mathcal{D}}(\pi)=\left(\prod_{s\in\mathcal{D}}\frac{C(s_{Oz})}{C(s_{t}^{\pi})}\right)^{\frac{1}{|\mathcal{D}|}}$ (8)
where $C$ counts the number of IR-instructions, $s_{Oz}$ is the IR optimized by LLVM -Oz flag, and $s_{t}^{\pi}$ is the final IR produced by the algorithm $\pi$. A value above $1$ indicates that the algorithm performs better than -Oz option on average.
Benchmarks
We evaluate our method mainly on benchmarks from the CompilerGym platform (Cummins et al. 2022): benchmark suites including cBench (Fursin et al. 2007), CHStone (Hara et al. 2008), MiBench (Guthaus et al. 2001), and NASA Parallel Benchmarks (NPB) (Bailey et al. 1995), in addition to kernels from open source projects including BLAS (Lawson et al. 1979), Linux, OpenCV (Culjak et al. 2012), and TensorFlow (Abadi et al. 2016). We disregard benchmarks generated by synthetic program generators (Lattner and Adve 2004; Yang et al. 2011) since they do not align with real-world scenarios. We follow the standard data splits of CompilerGym.
Additionally, we evaluate CompilerDream on a large-scale dataset named FormAI (Tihanyi et al. 2023). FormAI comprises a vast collection of AI-generated C programs with diverse functionalities and coding styles. We filtered out codes that failed to compile into CompilerGym benchmarks, resulting in a test set of 109,016 programs.
Inspired by FormAI, we also constructed a dataset of 50 Objective-C programs generated by a Large Language Model to further evaluate CompilerDream’s ability to generalize to different programming languages. All details on benchmarks can be found in Appendix C.
Figure 3: Generel code optimization results: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods. Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
4.2 Autotuning: CompilerGym Leaderboard
Method Walltime
Code size
Reduction
CompilerDream
   + Guided Search
60.8s 1.073$\times$
PPO + Guided Search 69.8s 1.070$\times$
CompilerDream 2.9s 1.068$\times$
Random Search ($t$=10800) 10,512.4s 1.062$\times$
Random Search ($t$=3600) 3,630.8s 1.061$\times$
Greedy Search 169.2s 1.055$\times$
GATv2 + DD-PPO 258.1s 1.047$\times$
Table 1: Autotuning results: CompilerDream’s code size reduction and time consumption on cBench dataset compared with other methods from CompilerGym leaderboard (Facebook 2022). Results for methods other than ours are taken directly from the leaderboard, where learning-based methods report only inference time as wall time.
We first use CompilerDream as an autotuning method to verify that its training process has sufficient capability to discover high-quality optimization sequences, supporting our goal of training an effective general code optimization agent.
Implementation
We target the task of CompilerGym leaderboard (Facebook 2022), optimizing pass sequences across 23 cBench programs. CompilerDream is trained for 25 hours, averaging about one hour per program—a time budget comparable to search methods. After training, besides evaluating single-trial optimization performance (more in Section 4.4), we test CompilerDream+Guided Search, following the leaderboard’s top method. This approach uses the agent’s policy $\pi_{\psi}(a_{t}|s_{t})$ to guide random search and limits search time to 1 minute per program (details in Appendix B.7).
Results
As shown in Table 1, CompilerDream’s policy achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench with a single trial after training, surpassing random search methods that require significantly more time—even when accounting for CompilerDream’s training time. This highlights the effectiveness of CompilerDream’s model-based RL training process in finding superior optimization passes, making it a promising component in building a general code optimizer. Moreover, when combined with guided search, CompilerDream outperforms the top method on the leaderboard while using less wall time.
4.3 General Value Prediction
We then demonstrate that the large-scale trained world model in CompilerDream can generalize to accurately simulate the optimization process of unseen input programs, thereby supporting the training of a general optimization agent. To illustrate this, we utilize our trained world model to estimate the outcome of applying an optimization sequence, a task named value prediction.
Implementation
We adopt the state-of-the-art Coreset-NVP method (Liang et al. 2023), which evaluates pass sequences within a fixed candidate set (the coreset with 625 sequence prefixes) and selects the best prefix via compiler validation. Coreset-NVP employs a neural network to directly predict the normalized value of sequences, which is learned via supervised learning. Our Coreset-CompilerDream method leverages a world model trained on CodeContests, to simulate the optimization process and accumulate predicted rewards as the value. Both our method and the baseline method Coreset-NVP are tested on four datasets orthogonal to their training sets. The number of compiler validation calls is set the same across methods to ensure a fair comparison. Additional implementation details are available in Appendix B.8.
Method Dataset
cBench CHStone MiBench NPB
Coreset-NVP 1.028 1.101 1.003 1.085
Coreset
-CompilerDream
1.038 1.101 1.017 1.140
Coreset-Oracle 1.041 1.106 1.020 1.159
Table 2: General value prediction results: The geometric mean of code size reduction achieved by the best sequences selected by different methods across 4 datasets. Results of Coreset-NVP and Coreset-Oracle is reported from Coreset-NVP’s original paper (Liang et al. 2023).
Results
Table 2 shows the geometric mean reduction achieved by different methods. Despite the test datasets being distinct from the training set, CompilerDream outperforms Coreset-NVP across three datasets, highlighting its strong compiler modeling and generalization capabilities. On MiBench, CompilerDream performs nearly optimally compared to the Oracle baseline, which uses a brute-force search and serves as the problem’s upper bound. Both our method and Coreset-NVP achieve an average reduction of $1.101$ on CHStone, close to the upper bound of $1.06$.
4.4 General Code Optimization
We finally aim to demonstrate that CompilerDream can learn a general optimization agent, capable of end-to-end generating effective optimization sequences for a wide variety of programs unseen during training. This scenario represents a typical daily use case where the optimization algorithm has limited time to compute the best solution for each of the diverse programs. Consequently, all methods generate only one optimization sequence per program in the test datasets unless otherwise specified.
Figure 4: Histograms of large-scale evaluations comparing CompilerDream and PPO on the FormAI dataset.
Figure 5: A comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model. The learned world model accurately captures the variations of program features and optimization metrics.
Figure 6: Analysis. Evaluations of different RL algorithms: (a) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set. A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends. (b) Zero-shot generalization capabilities of different RL algorithms on various test benchmark datasets. Effect of training dataset: (c) Test performance of CompilerDream trained on CodeContests and Csmith. Bars indicate the standard deviation.
Implementation
We compare CompilerDream with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method. The random search conducts hundreds of trials within a time budget similar to CompilerDream’s. The -Oz flag represents LLVM’s highest level of code size optimization while the -O0 flag represents no optimization. The learning-based method using Proximal Policy Optimization (PPO) (Schulman et al. 2017) and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs. Following the state-of-the-art approach, all methods operate within a limited action space.
Results
Figure 3 shows CompilerDream’s performance in reducing code size, measured by the geometric mean of IR instruction count reduction. Without specific in-domain training, CompilerDream outperforms -Oz in all but two test benchmarks in a single trial and surpasses PPO in all but the BLAS datasets. It also exceeds random search within a similar time budget on most datasets, except Linux. The minimal performance differences on the BLAS and Linux datasets suggest they are already highly optimized.
Moreover, CompilerDream’s zero-shot generalization matches or exceeds in-domain training. This advantage is especially notable in the NPB dataset, where data sparsity challenges in-domain agents, yet CompilerDream achieves an additional $3\%$ code size reduction. CompilerDream also generalizes well to new languages, evidenced in its performances on the BLAS and NPB datasets of Fortran programs. Further, on the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain instances.
On the large-scale FormAI test set (Figure 4), CompilerDream outperforms PPO, matching or outperforming -Oz on more programs and achieving higher optimization levels, demonstrating its superior and consistent performance.
4.5 Analysis
Comparison with model-free methods
We further evaluate the sample efficiency and zero-shot generalization abilities of CompilerDream based on the world model algorithm, Dreamer, against a range of model-free counterparts, including PPO (Schulman et al. 2017), DQN (Mnih et al. 2015), A2C (Mnih et al. 2016), APEX (Horgan et al. 2018), and IMPALA (Espeholt et al. 2018; Schulman et al. 2017). Figure 6a illustrates that while PPO is the strongest model-free baseline, CompilerDream outperforms it, learning an order of magnitude faster with fewer compiler interactions. As Figure 6b demonstrates, CompilerDream also excels in generalizing to unseen benchmarks, supporting our hypothesis that world model-based agents better capture the environment’s structure and enhance zero-shot generalization.
Effect of training dataset
To assess the impact of the CodeContests dataset on the generalization ability, we compared it with the commonly used Csmith (Yang et al. 2011) dataset, a large LLVM IR dataset generated by rules. We train CompilerDream on both datasets and the results shown in Figure 6 indicate that CompilerDream trained on CodeContests significantly outperforms the Csmith-trained version across all five test datasets. This advantage is particularly evident on the manually curated cBench (Fursin et al. 2007) and CHStone (Hara et al. 2008) datasets, demonstrating that the CodeContests-trained model can generalize more effectively to human-written programs.
Program showcase
In Figure 5, we display a predicted optimization trajectory for an unseen program from cBench, as forecasted by our learned compiler world model. The model successfully forecasts numeric features of future IR, including the counts of branches and blocks, alongside future rewards that signify optimization outcomes. This instance exemplifies the capability of our learned compiler world model to serve as a viable alternative for a real compiler environment in training code optimization agents.
5 Discussion
We aim to tackle the major challenge of zero-shot generalization that arises when applying learning-based code optimization. We thus introduce the CompilerDream approach that leverages the sample efficiency and generalization capabilities of world model-based reinforcement learning, as well as the strengths of a diverse, high-quality training program dataset. Our results have shown that CompilerDream, with its powerful generalization ability, attains optimization performance that outperforms existing methods, in multiple problem scenarios under the challenging area of phase ordering. There is substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world model, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models.
Appendix A Extended Related Work
Here we provide an overview of several topics in compiler optimization that our work did not focus on but are nonetheless important:
Feature Extraction
Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering. The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni, Dubach, and O’Boyle 2014; Haj-Ali et al. 2020), designed by expert intuitions. Numerous studies have aimed to reduce the cost of feature design. Following the success of word2vec embeddings in natural language processing (Mikolov et al. 2013), methods like code2vec (Alon et al. 2019), inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), and IR2vec (VenkataKeerthy et al. 2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR). The surge in deep learning has enabled feeding raw information such as AST (Alon et al. 2018), control-data flow graphs (CFG) (Brauckmann et al. 2020; Cummins et al. 2021), and code token sequences (Cummins et al. 2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
Large Language Model
With the rise of large language models (LLMs), recent work has introduced LLMs into compiler optimization problems, demonstrating remarkable effectiveness in predicting optimal optimization sequences and comprehending the optimization process (Cummins et al. 2023). LLMs can directly analyze the intermediate representation of code in textual form, enabling a better understanding of the code’s intricate details. However, due to limitations on input length, it has to split programs into shorter functions for training and prediction, preventing it from predicting on code of arbitrary length. We also want to emphasize that optimizing individual functions is quite different from optimizing an entire program. Many optimization passes, such as –inline and –mergefunc, which handle function inlining and merging, are either ineffective or less critical when there’s only one function to optimize.
Frameworks
Platforms that expose the compiler as a playground for AI experiments have significantly reduced the entry barriers to intelligent compiler research. OpenTuner (Ansel et al. 2014), and YaCoS (Filho, Rodriguez, and da Silva 2018) serve as autotuning frameworks with a range of compiler optimization techniques. Our experiments utilize CompilerGym (Cummins et al. 2022), which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner. We are optimistic that the future release of our compiler world model and code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
Appendix B Implementation Details
B.1 Compiler Environment
Our experiments are conducted on the CompilerGym platform (Cummins et al. 2022), version 0.2.5, with LLVM-10.0.0 integration.
B.2 Features and Actions
Index Name Description
0 BBNumArgsHi Number of BB where total args for phi nodes is gt 5
1 BBNumArgsLo Number of BB where total args for phi nodes is [1, 5]
2 onePred Number of basic blocks with 1 predecessor
3 onePredOneSuc Number of basic blocks with 1 predecessor and 1 successor
4 onePredTwoSuc Number of basic blocks with 1 predecessor and 2 successors
5 oneSuccessor Number of basic blocks with 1 successor
6 twoPred Number of basic blocks with 2 predecessors
7 twoPredOneSuc Number of basic blocks with 2 predecessors and 1 successor
8 twoEach Number of basic blocks with 2 predecessors and successors
9 twoSuccessor Number of basic blocks with 2 successors
10 morePreds Number of basic blocks with gt. 2 predecessors
11 BB03Phi Number of basic blocks with Phi node count in range (0, 3]
12 BBHiPhi Number of basic blocks with more than 3 Phi nodes
13 BBNoPhi Number of basic blocks with no Phi nodes
14 BeginPhi Number of Phi-nodes at beginning of BB
15 BranchCount Number of branches
16 returnInt Number of calls that return an int
17 CriticalCount Number of critical edges
18 NumEdges Number of edges
19 const32Bit Number of occurrences of 32-bit integer constants
20 const64Bit Number of occurrences of 64-bit integer constants
21 numConstZeroes Number of occurrences of constant 0
22 numConstOnes Number of occurrences of constant 1
23 UncondBranches Number of unconditional branches
24 binaryConstArg Binary operations with a constant operand
25 NumAShrInst Number of AShr instructions
26 NumAddInst Number of Add instructions
27 NumAllocaInst Number of Alloca instructions
28 NumAndInst Number of And instructions
29 BlockMid Number of basic blocks with instructions between [15, 500]
30 BlockLow Number of basic blocks with less than 15 instructions
31 NumBitCastInst Number of BitCast instructions
32 NumBrInst Number of Br instructions
33 NumCallInst Number of Call instructions
34 NumGetElementPtrInst Number of GetElementPtr instructions
35 NumICmpInst Number of ICmp instructions
36 NumLShrInst Number of LShr instructions
37 NumLoadInst Number of Load instructions
38 NumMulInst Number of Mul instructions
39 NumOrInst Number of Or instructions
40 NumPHIInst Number of PHI instructions
41 NumRetInst Number of Ret instructions
42 NumSExtInst Number of SExt instructions
43 NumSelectInst Number of Select instructions
44 NumShlInst Number of Shl instructions
45 NumStoreInst Number of Store instructions
46 NumSubInst Number of Sub instructions
47 NumTruncInst Number of Trunc instructions
48 NumXorInst Number of Xor instructions
49 NumZExtInst Number of ZExt instructions
50 TotalBlocks Number of basic blocks
51 TotalInsts Number of instructions (of all types)
52 TotalMemInst Number of memory instructions
53 TotalFuncs Number of non-external functions
54 ArgsPhi Total arguments to Phi nodes
55 testUnary Number of Unary operations
Table 3: Descriptions of 56-dimension Autophase features (adapted from (Cummins et al. 2022)).
We extract program features following the approach of Autophase (Haj-Ali et al. 2020). We use a 56-dimension Autophase feature vector concatenated with a 42-dimension action histogram vector as the observation. The Autophase feature contains various statistics of the IR code, with each dimension thoroughly explained in Table 3. The action histogram vector contains the counts of actions already taken by the agent within the current episode. Both vectors are normalized to ensure that their values fall within a reasonable range. Specifically, each element of the Autophase vector is divided by the total instruction count of the program, whereas the action histogram vector is normalized by the total action count per episode, set as 45.
We choose to use this feature extraction method because it is both efficient and effective. In our preliminary experiments, we find that more complex feature extraction methods, such as ProGraML (Cummins et al. 2021) and inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), significantly slow down the training and inference of CompilerDream, with only marginal improvements in performance and generalization.
We employ two distinct action spaces in our experiments to align with the baseline methods. In Sections 4.2 and Section 4.3, we utilize the full action space comprising all 124 optimization passes in LLVM. In Section 4.4 and Section 4.5, CompilerDream and the baseline methods are trained and tested on an action space derived from Autophase (Haj-Ali et al. 2020). The original version of this action space included 45 LLVM optimization passes. However, CompilerGym excludes 3 actions due to updates in the latest LLVM version. Thus, we utilize a total of 42 actions. Table 4 presents the names of the LLVM optimization passes corresponding to these 42 actions. Using this limited action space is a prevailing practice, and it has been proved to be effective in prior works(Haj-Ali et al. 2020; Cummins et al. 2022). It also makes the training easier since the model can focus on the most effective optimization passes.
Index Name Index Name Index Name
0 -adce 14 -instcombine 28 -lowerinvoke
1 -break-crit-edges 15 -ipsccp 29 -lowerswitch
2 -constmerge 16 -jump-threading 30 -mem2reg
3 -correlated-propagation 17 -lcssa 31 -memcpyopt
4 -deadargelim 18 -licm 32 -partial-inliner
5 -dse 19 -loop-deletion 33 -prune-eh
6 -early-cse 20 -loop-idiom 34 -reassociate
7 -functionattrs 21 -loop-reduce 35 -sccp
8 -functionattrs 22 -loop-rotate 36 -simplifycfg
9 -globaldce 23 -loop-simplify 37 -sink
10 -globalopt 24 -loop-unroll 38 -sroa
11 -gvn 25 -loop-unswitch 39 -strip
12 -indvars 26 -lower-expect 40 -strip-nondebug
13 -inline 27 -loweratomic 41 -tailcallelim
Table 4: A list of LLVM transformation passes selected in Autophase’s action space.
B.3 Reward Function
The reward function is defined as the normalized change of the optimization metric $C(s)$:
$r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$ (9)
where a lower $C$ indicates better performance. In line with other works focused on optimizing code size, we define $C(s)$ as the IR instruction count at compiler state $s$. The $C(s_{\mathrm{b}})$ stands for the performance of a baseline policy, specifically the built-in -Oz flag in our experiments. A total reward greater than 1 means that the optimization sequence performs better than the baseline policy.
B.4 Hyperparameters
The hyperparameters for our DreamerV3 agent implementation are outlined in Table 5. For hyperparameters not specified, we use the same value as the original DreamerV3 (Hafner et al. 2023). Most of the listed hyperparameter values are directly taken from DreamerV2 (Hafner et al. 2021) or previous works, with minimal tuning. However, the loss scales are carefully tuned, as the observation loss is relatively small compared to the reward loss in our setting, and the balance between these losses has a significant impact on CompilerDream’s performance. All experiments share the same set of hyperparameters unless otherwise specified.
Hyperparameter Value
Architecture RSSM recurrent units 1024
RSSM number of latents 32
RSSM classes per latent 32
MLP layers 4
MLP hidden units 400
Activation LayerNorm + SiLU
Training Random exploration 500 environment steps
Replay buffer capacity $2\times 10^{6}$
Reward smoothing $\alpha$ (Lee, Abbeel, and Lee 2024) 0.6
Training frequency Every 5 environment steps
Batch size 50
Batch length $T$ 50
Imagination horizon $H$ 15
Discount $\gamma$ 0.99
$\lambda$-target discount 0.95
World model loss scales 100.0 for Autophase
10.0 for action histogram
1.0 for reward
5.0 for discount
0.1 for KL
Actor entropy regularization $\eta$ $3\times 10^{-4}$
KL balancing 0.8
Optimizer Adam
World model learning rate $1\times 10^{-4}$
Actor-critic learning rate $3\times 10^{-5}$
Weight decay $1\times 10^{-5}$
Gradient clipping 100
Table 5: Hyperparameters of DreamerV3 in our experiments.
B.5 Training Model-free Baselines
We use RLlib (Liang et al. 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al. 2017) (Haj-Ali et al. 2020), A2C (Mnih et al. 2016), IMPALA (Espeholt et al. 2018), APEX (Horgan et al. 2018), and DQN (Mnih et al. 2015). All the experiments are conducted with 10 CPUs and an RTX-3090 GPU. For the PPO baseline in Section 4.4, we trained it for around 7 hours since we observed overfit to CodeContests when training for a longer time. Other model-free methods are trained for at least 10 hours. The training process uses 5 workers to interact with the environment and use 4 evaluation workers to evaluate the checkpoint on the validation set. We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al. 2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 6. We explored roughly 2 or 3 values for each hyperparameter listed and selected the set of hyperparameters that yielded the best performance on the validation set. The results of the PPO baseline in Section 4.5 and Figure 6 are exactly the same as those in Section 4.4 and Figure 3.
Hyperparameters Value
PPO gamma 1.0
use_gae True
lambda_ 1.0
train_batch_size 9000
lr 5e-5
kl_coeff 0.2
kl_target 0.01
vf_loss_coeff 1.0
num_sgd_iter 30
sgd_minibatch_size 128
clip_param 0.3
vf_clip_param 10.0
weight_decay 1e-6
Table 6: Hyperparameters for the PPO baseline, well-tuned on our dataset to be deviating from the default value in RLlib.
B.6 Autotuning
In Section 4.2, we extensively trained CompilerDream on the cBench dataset to evaluate its effectiveness as an autotuning method in comparison to other approaches on CompilerGym’s leaderboard. Similar to other studies, we trained CompilerDream using all programs in cBench, excluding the one called ghostscript due to its considerable size, which would significantly hinder the training process. The action space is set to the full set of 124 actions in LLVM. All other training details remained unchanged.
B.7 Guided Search
The guided search strategy we use in Section 4.2 followed the design in the PPO+Guided Search baseline. During the search, actions are sampled from the agent’s learned policy $\pi_{\psi}(a_{t}|s_{t})$ up to 45 steps, at which point a new episode begins. Code size reductions after applying each action are all recorded to calculate the maximum. To enhance exploration, we add a uniform distribution to the policy, implemented by a $5\%$ chance of selecting a uniformly random action instead of sampling from the policy. At each step, the code size reduction achieved so far is recorded for later comparison, and the elapsed time is monitored to ensure that the search is completed within the 1-minute time limit per benchmark. Compared to the guided search in PPO+Guided Search (in Table 1), we use a horizon of 45 steps instead of 200 steps. Additionally, PPO+Guided Search includes an extra search of 500 steps on the best sequence found, which we do not include, resulting in our wall time being a few seconds shorter than theirs.
B.8 Value Estimation
Here, we provide additional details about the experiment in Section 4.3. The baseline method, Coreset-NVP (Liang et al. 2023), trains a value prediction model that scores sequences from a fixed set of optimization sequences, referred to as the core set, which is obtained through an extensive search on their training set. For each program to optimize, it enumerates the first 45 passes in the top-scored sequences and validates their effectiveness through actual compilation. Specifically, sequences are processed in descending order of scores, executing each pass within a sequence. The compiler state (i.e., the IR) resets to its initial state after all passes in a sequence have been enumerated, before moving on to the next sequence. The code size reduction achieved after each pass is recorded, and the pass, along with its corresponding sequence prefix that yields the highest code size reduction, is selected. The core set in Coreset-NVP contains 50 sequences comprising a total of 625 passes, resulting in an average sequence length of 12.5 passes. Thus, roughly 3 or 4 sequences are attempted per benchmark, given the limit of 45 passes. These passes correspond to the optimization actions defined in Section 3.1, with a total of 79 different optimization passes included.
For our experiment, we set the action space to the full set of all 124 optimization passes available in LLVM when training our world model. Focusing on predicting the target value of cumulative reward after executing each optimization pass in the core set, we replace CompilerDream’s optimization agent with a policy that sequentially outputs the sequences from the core set, allowing the world model to focus on the problem dynamics driven by these sequences. This setup is also comparable to the baseline, as Coreset-NVP employs a value predictor that outputs a 50-dimension vector corresponding to the normalized value of each sequence in the core set. That prediction model is obtained through supervised training.
We train our world model using the CodeContests dataset introduced in Section 3.4, consistent with the experiments in Section 4.4. During training, we increase the weight of the reward loss to 100.0, while keeping the other hyperparameters unchanged. To align with the methodology used in Coreset-NVP, we validate the top 45 sequence prefixes of each program with the highest predicted values through actual compilation and select the best-performing sequence. The world model is trained using three different random seeds, and the average results are reported in Table 2.
B.9 Hardware and Training Time
We train all CompilerDream-based methods with 64 CPUs and an RTX-3090 GPU. In Section 4.4, we trained CompilerDream on the CodeContests dataset for around 1 day and 20 hours. In Sections 4.2 and 4.3, CompilerDream was trained for about 1 day.
The random search baseline in Figure 3 is conducted with 4 CPUs, which is sufficient since it is a single-thread program.
The random search baseline in Table 1 utilizes 80 CPUs, as specified in the write-up attached to CompilerGym’s leaderboard (Facebook 2022). Our CompilerDream + Guided Search method in Table 1 is tested on the same machine used for training CompilerDream, equipped with 64 CPUs and an RTX-3090 GPU.
We ensure that sufficient memory is available for all experiments. Memory size should have little impact on the results in all of our experiments.
B.10 Random Seeds
All of our experiments that report a min-max range or standard deviation were replicated using three different random seeds. These seeds were randomly selected, and the results are generally reproducible with other seeds as well.
Appendix C Benchmarks
CompilerGym Benchmarks
In our study, we select eight benchmarks for zero-shot test and in-domain training: BLAS, cBench, CHStone, Linux, MiBench, NPB, OpenCV, and TensorFlow. These benchmarks are part of the built-in datasets provided by CompilerGym version 0.2.5. For benchmarks with a total number of programs more than 100, we use the first 50 programs as the test set, the following 50 programs as the validation set, and all of the rest as the training set. These training and validation sets are only used for in-domain training. The datasets comprising fewer than 100 programs are not applicable for in-domain training; instead, all their programs are allocated to the test set. The number of programs in each dataset after division is detailed in Table 7.
Dataset Training Split Validation Split Test Split
BLAS 200 50 50
cBench 23 N/A N/A
CHStone 12 N/A N/A
Linux 13,794 50 50
MiBench 40 N/A N/A
NPB 22 50 50
OpenCV 342 50 50
TensorFlow 1,885 50 50
Table 7: Dataset division of 8 CompilerGym benchmarks.
AI-Generated Benchmarks
To further test the generalization ability of our CompilerDream agent on different programming languages, we borrow the method from FormAI (Tihanyi et al. 2023) and generate a dataset containing 50 unique Objective-C programs using GPT-3.5. We use the same prompt as FormAI, except that we add an instruction to ask GPT to generate programs that can be directly compiled under Clang version 10.0.0 and do not use ARC (Automatic Reference Counting), to improve the compilation pass rate of generated programs. We compile the generated programs using Clang without including any third-party libraries, and all programs that cannot pass compilation are discarded.
Appendix D Extended Experimental Results
Dataset Program Pass Sequences by CompilerDream Agent Code size
O0 CompilerDream Oz (Reduction)
cBench sha -sroa -gvn -sroa -sroa -instcombine -simplifycfg -sroa -functionattrs -licm -functionattrs -early-cse -simplifycfg -indvars -gvn -simplifycfg -early-cse -functionattrs -reassociate -memcpyopt -gvn -simplifycfg -reassociate -functionattrs -early-cse -instcombine -early-cse -lowerinvoke -simplifycfg -functionattrs -functionattrs -memcpyopt -gvn -simplifycfg -functionattrs -reassociate -early-cse -instcombine -early-cse -strip -strip -early-cse -strip -early-cse -strip -early-cse 799 349 500 (1.43$\times$)
bzip2 -sroa -gvn -sroa -sroa -simplifycfg -strip -strip -instcombine -simplifycfg -early-cse -simplifycfg -strip -strip -reassociate -reassociate -memcpyopt -jump-threading -memcpyopt -functionattrs -memcpyopt -functionattrs -licm -functionattrs -gvn -simplifycfg -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -jump-threading -memcpyopt -early-cse -strip -strip -instcombine -simplifycfg -jump-threading -jump-threading -early-cse -jump-threading -early-cse -simplifycfg -strip -functionattrs 28748 13565 15946 (1.18$\times$)
OpenCV #41 -sroa -early-cse -early-cse -simplifycfg -early-cse -early-cse -early-cse -sink -sink -sink -loop-deletion -loop-deletion -loop-deletion -lowerinvoke -simplifycfg -sroa -sroa -sroa -early-cse -early-cse -early-cse -early-cse -lowerinvoke -lowerinvoke -lowerinvoke -globalopt -sroa -instcombine -simplifycfg -early-cse -early-cse -early-cse -sink -globalopt -sroa -sroa -functionattrs -simplifycfg -early-cse -functionattrs -early-cse -sink -functionattrs -functionattrs -functionattrs 28 18 28 (1.56$\times$)
#9 -sroa -gvn -sroa -early-cse -simplifycfg -strip -strip -loop-deletion -lowerinvoke -simplifycfg -sroa -early-cse -instcombine -simplifycfg -early-cse -memcpyopt -early-cse -strip -strip -strip -functionattrs -instcombine -simplifycfg -jump-threading -simplifycfg -memcpyopt -early-cse -simplifycfg -reassociate -memcpyopt -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -instcombine -simplifycfg -jump-threading -early-cse -jump-threading -strip -simplifycfg -early-cse -simplifycfg -functionattrs 9510 6341 9269 (1.18$\times$)
TensorFlow #17 -sroa -gvn -early-cse -simplifycfg -early-cse -early-cse -instcombine -early-cse -sink -sink -sink -lowerinvoke -simplifycfg -sroa -early-cse -memcpyopt -early-cse -memcpyopt -instcombine -simplifycfg -early-cse -early-cse -early-cse -simplifycfg -early-cse -early-cse -strip -strip -indvars -simplifycfg -early-cse -memcpyopt -early-cse -instcombine -early-cse -early-cse -simplifycfg -early-cse -memcpyopt -early-cse -early-cse -instcombine -simplifycfg -early-cse -early-cse 5512 4247 5450 (1.28$\times$)
Table 8: Top performances of the zero-shot CompilerDream agent on individual programs. Full IRs of these programs before and after optimizations are shown in the Supplementary Information.
D.1 Learning Curves
The zero-shot test performance of our CompilerDream agents during training is shown in Figure 7. Note that we validate and test the agent every 10000 environment steps and report the test performance from the checkpoint that achieved the best validation results for comparison among various methods.
Figure 7: Zero-shot test performance of the CompilerDream agent during training. We report mean and standard deviation across three runs.
D.2 Quantitative Results
Quantitative results corresponding to Figure 3 in the main text are provided in Table 9.
D.3 Program Case Study
In Addition to the predicted optimization trajectory showcase in our paper, we present more example of optimization results here. Table 8 presents the agent’s top performance outcomes on various benchmark datasets. We observe that the agent indeed produces a specialized optimization strategy tailored for each program. Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -early-cse(early common subexpression elimination) and -simplifycfg, as particularly effective in code size optimization.
Figure 8: t-SNE (Van der Maaten and Hinton 2008) visualization of programs from training and test datasets.
D.4 Data Distribution Visualization
In Figure 8, we visualize the distribution of our training and test datasets. To accurately represent the dynamic behavior of programs, we randomly select 1000 action sequences, each with a length of 45, from our action space. These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program. These comprehensive feature vectors are finally processed using t-SNE (Van der Maaten and Hinton 2008) for dimensionality reduction and visualization.
Figure 8 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles). This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments (Cummins et al. 2022), which shows a significant deviation from real-world applications. Nonetheless, our visualization can still not perfectly capture the transferability across datasets. For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.
-O0 Random Policy
Random Search
Autophase
CompilerDream
(In-domain)
CompilerDream
(Zero-shot)
BLAS geomean 0.931 0.821$\pm$0.018 0.960$\pm$0.004 0.993$\pm$0.002 0.993$\pm$0.005 0.991$\pm$0.000
min 0.707 0.432$\pm$0.106 0.723$\pm$0.076 0.916$\pm$0.004 0.831$\pm$0.148 0.913$\pm$0.000
max 1.000 1.002$\pm$0.004 1.031$\pm$0.010 1.014$\pm$0.001 1.025$\pm$0.005 1.016$\pm$0.000
cBench geomean 0.481 0.737$\pm$0.015 0.858$\pm$0.020 1.010$\pm$0.004 N/A 1.036$\pm$0.002
min 0.299 0.401$\pm$0.097 0.560$\pm$0.030 0.817$\pm$0.109 N/A 0.962$\pm$0.001
max 0.626 1.150$\pm$0.113 1.298$\pm$0.083 1.405$\pm$0.014 N/A 1.395$\pm$0.054
CHStone geomean 0.487 0.701$\pm$0.043 1.037$\pm$0.006 1.071$\pm$0.002 N/A 1.094$\pm$0.005
min 0.402 0.278$\pm$0.108 0.889$\pm$0.044 0.967$\pm$0.027 N/A 1.000$\pm$0.006
max 0.655 1.038$\pm$0.033 1.304$\pm$0.038 1.313$\pm$0.045 N/A 1.378$\pm$0.021
Linux geomean 0.988 0.981$\pm$0.004 1.001$\pm$0.000 0.985$\pm$0.001 0.993$\pm$0.000 0.986$\pm$0.005
min 0.615 0.740$\pm$0.126 1.000$\pm$0.000 0.615$\pm$0.000 0.700$\pm$0.001 0.642$\pm$0.041
max 1.011 1.001$\pm$0.002 1.011$\pm$0.000 1.009$\pm$0.003 1.011$\pm$0.000 1.009$\pm$0.003
Mibench geomean 0.389 0.812$\pm$0.022 1.005$\pm$0.001 1.000$\pm$0.007 N/A 1.006$\pm$0.002
min 0.278 0.352$\pm$0.074 0.858$\pm$0.010 0.833$\pm$0.062 N/A 0.879$\pm$0.003
max 0.760 1.392$\pm$0.074 1.603$\pm$0.015 1.524$\pm$0.045 N/A 1.588$\pm$0.000
NPB geomean 0.530 0.833$\pm$0.011 1.074$\pm$0.010 1.071$\pm$0.018 1.075$\pm$0.017 1.108$\pm$0.001
min 0.191 0.414$\pm$0.043 0.805$\pm$0.104 0.801$\pm$0.070 0.810$\pm$0.082 0.886$\pm$0.000
max 1.066 1.848$\pm$0.193 2.141$\pm$0.062 2.235$\pm$0.101 2.315$\pm$0.028 2.343$\pm$0.012
OpenCV geomean 0.981 0.949$\pm$0.011 1.080$\pm$0.001 1.082$\pm$0.004 1.087$\pm$0.007 1.092$\pm$0.000
min 0.833 0.603$\pm$0.133 0.888$\pm$0.021 0.897$\pm$0.002 0.898$\pm$0.001 0.897$\pm$0.000
max 1.370 1.409$\pm$0.167 1.571$\pm$0.022 1.620$\pm$0.193 1.635$\pm$0.186 1.556$\pm$0.000
TensorFlow geomean 0.983 0.912$\pm$0.009 1.006$\pm$0.003 1.010$\pm$0.019 1.032$\pm$0.000 1.032$\pm$0.001
min 0.927 0.618$\pm$0.046 0.877$\pm$0.009 0.841±0plus-or-minus0$\pm$ 0± 0.134 0.970$\pm$0.010 0.968$\pm$0.011
max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.281$\pm$0.006 1.289$\pm$0.008 1.282$\pm$0.002
Table 9: Quantitative results for code size reduction, corresponding to Figure 3. We report mean and standard deviation across three runs.